# Better Together? A Field Experiment on

### v2

@@ 5.2 Improved Targeting Reduced Child Harm and Disparities
Notes: This figure shows that groups that experienced the greatest improvements in the targeting of screen-ins also experienced the greatest reductions in overall child harm. The thick, dark red lines are normalized control group outcomes in the absence of any improvement from algorithm support. In contrast, the dashed purple lines represent average realized effects of providing the algorithmic tool on the harm index (vertical axis) and targeting (horizontal axis). Each dot represents the average effects of providing the algorithmic tool on harm and targeting for each group, respectively. The horizontal axis measures targeting by leveraging the difference in harm after the referral among screened-out children with the tool compared to screened-out children without the tool (relying on random assignment to treatment and control, which implies that harm would have been the same without algorithm-induced improvements in allocation). This corresponds to the targeting test reported in Column 1 of Table 3.
To further probe that improved screen-in targeting is a central mechanism for improvements in child outcomes, we conduct an additional test inspired by Figure 3. We randomly draw a thousand subsamples of varying size, mimicking sample size heterogeneity in the prior figure, and imposing a minimum sample size of $n=100$. For each random subsample, we then calculate the effect of the algorithm on the overall harm index and on targeting of screen-ins (the respective vertical and horizontal axes of Figure 3). Results are presented in Appendix Figure A.13. On average, a 1 SD improvement in targeting for screen-outs (corresponding to a 2 SD improvement in targeting for screen-ins, given the 2:1 screen-out to screen-in ratio) is associated with a reduction in overall child harm by 0.66 SD ($p<0.001$). The groups that benefit most from the algorithmic tool are the ones where targeting of screen-ins improves the most. This result is generalizable and not specific to the groups we report in Figure 3. Consistent with novel ongoing work by Lacey et al. (2024), we find that screening in a family for investigation and possible services can improve child health and wellbeing.
@@ 5.3 Worker Discussion Text: Changes in Worker Attention
How were workers able to better target screen-ins? In this section we provide suggestive evidence of how access to algorithmic information changed workers’ decision-making process. To do so, we leverage rich, unstructured administrative text data from the partnered child welfare agency, taken from discussion notes written during workers’ decision-making process (see examples in Table LABEL:table_reasons).
Combining these data with recent advances in large language models (LLMs), we assessed in a principled way whether workers discussions changed when they had access to the tool. Inspired by Ludwig et al. (2017) and using recent improvements in open-source LLM methods, we fine-tuned a pretrained LLM, DistilBERT, to predict whether workers had access to algorithm support using the notes from their discussions. [^15] We thank an anonymous referee for this suggestion. The model provides clear evidence that discussions changed as shown in Figure A.14 ($AUC\approx 0.7$).
Given the notorious difficulty in explaining “black-box” LLM predictions, we grouped together categories of words in light of the ways in which the algorithmic tool could modify attention. The algorithmic tool was based on data from household members’ past history and listed a score for each child on the report. With algorithm support, workers might make fewer mistakes in aggregating child or family history, better understand family history, better contextualize the incoming allegation report, or have a broader household focus that considers the circumstances of all siblings and alleged perpetrators. Each of these interrelated potential mechanisms could improve worker assessment of risk.
@@ 5.4 Human-plus-Algorithm Gains under Complexity
Algorithm Available $\cdot$ Type -0.153*** -0.106* -0.362**
@@
Frequency of Type 0.519 0.266 0.156
Observations 3,431 2,741 3,431
@@
Consistent with theory, gains from algorithm support were greatest among complex referrals. Across all three complexity measures, reductions in harm from the algorithmic tool were almost exclusively concentrated among complex referrals (Table 4, compare rows “Algorithm Available $\cdot$ Type” and “Algorithm Available”). The findings also hold when controlling for algorithm risk score. Although each of the three complexity measures are correlated with above-average levels of harm in the control group (Table 4, row “Type”), these findings are consistent with recent work showing that algorithm support is particularly beneficial to individuals for whom a task is complex (Brynjolfsson et al., 2023).
Did the algorithmic tool’s informational content (risk score) matter, or did its presence merely serve as a behavioral reminder or nudge to workers? Although difficult to answer definitively, the fact that workers regularly checked and recorded the algorithm score – or lack thereof – in both the intervention and control groups (Appendix Figure A.5), together with score-based heterogeneity in discussion content and decisions (Figures 2 and 4) suggest that benefits of the algorithm are tied to the process of reviewing the predicted risk scores. Although the specific 1-to-20 value of the algorithm’s risk score likely helped workers allocate their attention more effectively, we cannot definitively rule out benefits from a general increase in the salience of family history – or other characteristics – upon seeing a listed score, regardless of its magnitude. An important topic for future research is understanding the impacts of how information is presented to users in similar settings.
@@
To what extent do human decision-makers contribute their independent expertise to improve upon algorithmic predictions? As machine learning decision supports become more widespread, so do questions about human experts’ contributions to decisions, and when to defer to machine decisions. In this study, outcomes improved when workers were given access to an algorithm. Workers, however, did not target investigations toward children with higher algorithm risk scores, but instead used complementary information to target more effectively on other dimensions. Given how well the algorithm risk score predicts relevant outcomes (Appendix Figure A.3), it is reasonable to ask: Could outcomes have improved even more if workers had based decisions more closely on the algorithm risk score? And would there have been implications for equity?
@@ 6.1 Counterfactual 1: Deferring Entirely to an Algorithm
The first counterfactual exercise estimates the impact of fully deferring screening decisions to the algorithm using a bounding approach. The experiment did not include an algorithm-only treatment arm due to legal constraints: the statutory requirements for a screen-in do not always correspond to high algorithm-predicted risk.
By design, the algorithm provided a prediction of child risk, rather than an explicit recommendation of whether to screen in. For the counterfactual estimation, we follow a simple decision rule where the algorithm screens in children with the highest predicted risk. The counterfactual exercise operates as follows:
@@
Results of this counterfactual exercise are plotted in Figure 5. We compare child harm under a best case for our algorithm-only scenario (dashed orange line) with outcomes observed under human-only (dashed light blue line) and human-plus-algorithm decision-making (solid blue line with confidence band). The algorithm-only counterfactual outperforms the human-only scenario by construction, because it was assumed that children screened in by humans but screened out by the algorithm were not harmed by the screen out. This assumption is visualized by the algorithm-only line falling below the human-only line. The horizontal axis varies the effect of a screen-in for children that humans screened out but the algorithm would have screened in. These hypothetical screen-in effects range from zero to four standard-deviation reductions in harm. The vertical axis shows the average child harm index across the sample, for each hypothetical effect of a screen-in. We expect the actual treatment effect of a screen-in to be toward the far left of the axis (<1<1 SD reduction in harm), given that a screen-in is not a panacea, and given that the gap in harm between the most severely maltreated children (those in foster care) and the remainder of the sample is less than 0.5 standard deviations. Therefore, under reasonable assumptions that favor the algorithm, we find that child outcomes under a human-plus-algorithm scenario are better than, though statistically indistinguishable from, the algorithm-only best-case rule under realistic screen-in treatment effects (<1<1 SD), and similar under larger treatment effects. This result highlights the limits to fully automating complex social decisions with an algorithm, and provides evidence in favor of human-algorithm complementarity in our setting.
This counterfactual approach allows a researcher to estimate the impact of alternative screen-in decision rules, including new hypothetical algorithms not included in the experiment. For example, we find that a screen-in decision rule based on predicted harm index is similarly unlikely to outperform observed human-plus-algorithm outcomes (not shown), though possibly owing to low prediction power (low $R^{2}$). We also estimate a limit on how much child harm could be reduced under a perfect decision protocol (i.e., an ideal algorithm), conditional on having the same screen-in rate. Appendix Figure A.15 presents outcomes under an “oracle” screen-in decision rule where children are ranked by ex-post realized harm. According to this decision rule, the 30 percent of children with the highest ex-post realized harm index are selected to be screened in. This is an impractical decision rule because potential future harm is uncertain at the time of screen-in, so the estimate provides a best-case scenario for potential improvements in child wellbeing. Appendix Figure A.15 shows that this unrealistic oracle bound outperforms humans with the algorithm if assuming large benefits of a screen-in (>0.5>0.5 SD reduction in harm). As before, this bound assumes that children who humans screened in, but who the new decision rule would have screened-out, would not be harmed from a screen-out (two thirds of screened-in families in our setting would have been screened out with this new decision rule). The oracle exercise highlights that there could exist algorithms that outperform the intervention arm of our trial, perhaps using richer data sources or optimized for human use (Ludwig and Mullainathan, 2024).
@@ 6.2 Counterfactual 2: Marginal Increased Reliance on Algorithm
In contrast to the first counterfactual exercise, which deferred decisions entirely to the algorithm, the second counterfactual exercise examines whether social workers should have relied marginally more on the algorithm’s risk score to make screen-in decisions. The algorithm used only historical data, whereas social workers had access to both historical data and real-time, detailed caller reports about why a family was being referred. Since the algorithm was designed to complement human decision-making, a relevant counterfactual exercise is to evaluate a marginal increase in using the algorithm score to guide decisions, rather than fully deferring to the algorithm decision rule. Increased marginal reliance corresponds to screening in additional children that the algorithm predicted as high risk. A real-life analog of this would be a mandatory screen-in protocol for families scoring above a high-risk threshold (Lacey et al., 2024).
@@
We find that potential benefits from deferring more to an algorithm have already been realized by standard human oversight over algorithm predictions. Figure 6 shows realized child harm for children not investigated by CPS, by treatment status and algorithm-predicted risk level. Each point represents the mean harm index of screened-out children with a given algorithm risk score. Among these children, an algorithm-only decision-maker would first prioritize screening in those with the highest risk scores, on the right side of the figure. We first note a striking opportunity for potential welfare improvements under algorithmic decision support: when the tool was not available, children with a score of 20 who were not investigated experienced harm levels almost 0.6 SD above the mean (light blue line with circles). Social workers were missing out on a subset of high-scoring children in the control group who accrued significant harm, such that a mandatory screen-in protocol for these children might have readily improved outcomes. In contrast, worker access to algorithmic information (dark blue line with diamonds) improved screen-in targeting and greatly attenuated any potential benefits from mandating screen-ins for top risk scores, despite the fact that screen-out rates remained unchanged at the top of the score distribution. With algorithm access, the reduction in harm for screened-out children on referrals scoring a 20 is statistically significant at the 10 percent level ($\beta=-0.44,p=0.071$). For high algorithm risk scores, the mean harm for screened-out children under algorithm availability is approximately 0 to 0.1 SD greater than the control group mean. In other words, most of the capacity for harm reduction is already realized by social workers using the algorithm freely, such that further gains from a screen-in mandate would be comparatively modest, especially relative to the costs of additional investigations that would result from such a policy. Requiring a mandatory screen-in for all children with an algorithm score of 20 would result in a sizable 20 percent (6 percentage point) increase in overall screen-in rates. [^16] Due to a temporary imperfection from the implementing partner in setting probabilistic score cutoffs, slightly more than five percent of children received an algorithm score of 20 during the trial period.
@@
Figure 6 therefore highlights two key points about human-plus-algorithm gains. First, the algorithm helped workers target investigations more effectively. Although workers did not change their screen-in rates for children with the highest algorithm risk scores when using the tool (Figure A.11), workers did screen in a set of ex-post more vulnerable children. Second, workers used complementary information not available to the algorithm to reduce harm. In situations where workers “disagreed” with the algorithm (high algorithm score, worker screen-out), child outcomes were not substantially worse than those for other screened-out children in the sample. Therefore, ex-post partial automation of investigation decisions using algorithm predictions alone is likely to be of limited benefit in our setting. Future work with larger study samples could consider partial automation with optimal triage, as proposed by Raghu et al. (2019).
@@ 6.3 Counterfactual 3: Algorithms and Disparities without Human Oversight
Finally, we estimate how an algorithm-only decision rule would have impacted equity. This can be done using the initial counterfactual exercise described in Section 6.1. Reassuringly, the algorithm-only counterfactual for screen-in disparities is nearly assumption-free: Hypothetical screen-in recommendations can be directly deduced from ordering algorithm risk scores from high to low (Section 6.1). We do not need to make any assumptions about how a screen-in impacts child harm. Figure 7 shows CPS surveillance (screen-in) disparities for relevant demographic subgroups under three decision regimes: human-only, human-plus-algorithm, and algorithm-only.
@@
Notes: This figure shows point estimates and 90 percent confidence intervals for disparities in screen-in rates by demographic category. Human-only and human-plus-algorithm disparities are directly observed in the data. Algorithm-only disparities are inferred with minimal assumptions: control group observations are sorted by algorithm risk score, and mandatory screen-ins are imposed above a score threshold that keeps overall screen-in rates constant. The vertical axis shows the disparity in screen-in rates. A zero point estimate represents zero disparity in screen-ins and is illustrated using a horizontal dashed black line. For example, in the first panel (“Black vs. Other”) when the humans do not have access to the algorithm (human only, light blue triangle on the left) screen-in rates would be 20 percentage points greater for Black children compared to other children. Low income refers to children whose families are receiving SNAP benefits.
We find that human-plus-algorithm decisions improved equity not only compared to the status quo of human experts alone, but also compared to counterfactual screen-ins from the algorithm without human oversight. Relative to an algorithm-only decision rule, human oversight reduced disparities in CPS investigations by race, ethnicity, gender, and socioeconomic status: the dark blue squares in Figure 7 (human with algorithm) are always closer to zero than the orange diamonds (algorithm only), though point estimates are in some cases imprecise. Second, the algorithm on its own would have investigated significantly more Hispanic and low-income children relative to the status quo of human experts on their own (light blue triangles). This second finding is to some extent consistent with policy makers’ concerns that algorithmic tools could perpetuate historical bias in CPS intervention. Notably, increased surveillance under the algorithm-only counterfactual does not correspond with a reduction in harm for these subgroups relative to observed human-plus-algorithm decision-making (Appendix Figure A.16), suggesting that the algorithm-only disparities may not be warranted. In short, the improvements in equity observed during the trial would not have been fully realized – and for some groups may have worsened – without human oversight of the algorithm. Providing algorithm support to humans leads to more equitable levels of surveillance by CPS, relative to the algorithm on its own.

### v3

@@ 5.2 Improved Targeting Reduced Child Harm and Disparities
To further probe that improved screen-in targeting is a central mechanism for improvements in child outcomes, we conduct an additional test inspired by Figure 3. We randomly draw a thousand subsamples of varying size, mimicking sample size heterogeneity in the prior figure, and imposing a minimum sample size of $n=100$. For each random subsample, we then calculate the effect of the algorithm on the overall harm index and on targeting of screen-ins (the respective vertical and horizontal axes of Figure 3). Results are presented in Appendix Figure A.14. On average, a 1 SD improvement in targeting for screen-outs (corresponding to roughly a 2 SD improvement in targeting for screen-ins, given the 2:1 screen-out to screen-in ratio) is associated with a reduction in overall child harm by 0.72 SD ($p<0.001$). The groups that benefit most from the algorithmic tool are the ones where targeting of screen-ins improves the most. This result is generalizable and not specific to the groups we report in Figure 3. Consistent with novel ongoing work by Lacey et al. (2024), we find that screening in a family for investigation and possible services can improve child health and wellbeing.
@@
Improvements in Outcomes are Tied to Workers’ Screen-In Allocation Decisions
Notes: This figure shows that groups that experienced the greatest improvements in the targeting of screen-ins also experienced the greatest reductions in overall child harm. The thick, light blue lines are normalized control group outcomes in the absence of any improvement from algorithm support. In contrast, the dashed dark blue lines represent average realized effects of providing the algorithmic tool on the harm index (vertical axis) and targeting (horizontal axis). Each dot represents the average effects of providing the algorithmic tool on harm and targeting for each group, respectively. The horizontal axis measures targeting by leveraging the difference in harm after the referral among screened-out children with the tool compared to screened-out children without the tool (relying on random assignment to treatment and control, which implies that harm would have been the same without algorithm-induced improvements in allocation). This corresponds to the targeting test reported in Column 1 of Table 2.
@@ 5.3 Worker Discussion Text: Changes in Worker Attention
How were workers able to better target screen-ins? In this section we provide suggestive evidence of how access to algorithmic information changed workers’ decision-making process. To do so, we leverage rich, unstructured administrative text data from the partnered child welfare agency, taken from discussion notes written during workers’ decision-making process (see examples in Table LABEL:table_reasons). Discussion notes vary considerably in detail and quality, but may still provide valuable insight into whether discussions changed.
Combining these data with recent advances using word embeddings from sentence transformers, we assessed in a principled way whether workers’ discussions changed when they had access to the tool. Specifically, we trained a BERTopic model (Grootendorst, 2022) on discussion notes from approximately 11,000 referrals. We then used the generated topics ($t=131$) to predict whether workers had access to the algorithm tool in our trial sample, using three-fold cross-validation. Even with a drastically reduced dimensionality of the text data to 131 topics, the model provides clear evidence that discussions changed when workers had access to the algorithm, as shown in Figure A.15 ($AUC\approx 0.55$). A linear regression of treatment status on the selected topics gives an F-statistic of $F=77,p<0.0001$ (not reported).
Topics types found by the BERTopic model are not designed to predict what is changing in the discussions and still suffer from lack of explainability. For those reasons, we grouped together categories of words in light of the ways in which the algorithmic tool could modify attention. The algorithmic tool was based on data from household members’ past history and listed a score for each child on the report. With algorithm support, workers might make fewer mistakes in aggregating child or family history, better understand family history, better contextualize the incoming allegation report, or have a broader household focus that considers the circumstances of all siblings and alleged perpetrators. Each of these interrelated potential mechanisms could improve worker assessment of risk.
@@ 5.4 Human-plus-Algorithm Gains under Complexity
Algorithm Available $\cdot$ Type -0.097** -0.098* -0.314**
@@
Frequency of Type 0.519 0.266 0.157
Observations 4,681 3,754 4,681
@@
Consistent with theory, gains from algorithm support were greatest among complex referrals. Across all three complexity measures, reductions in harm from the algorithmic tool were almost exclusively concentrated among complex referrals (Table 3, see rows “Algorithm Available $\cdot$ Type” and “Algorithm Available”). The findings also hold when controlling for algorithm risk score. Although each of the three complexity measures are correlated with above-average levels of harm in the control group (Table 3, row “Type”), these findings are consistent with recent work showing that algorithm support is particularly beneficial to individuals for whom a task is complex (Brynjolfsson et al., 2023).
Did the algorithmic tool’s informational content (risk score) matter, or did its presence merely serve as a behavioral reminder or nudge to workers? Although difficult to answer definitively, the fact that workers regularly checked and recorded the algorithm score – or lack thereof – in both the intervention and control groups (Appendix Figure A.4), together with score-based heterogeneity in discussion content and decisions (Figures 2 and 4) suggest that benefits of the algorithm are at least partially attributable to the process of reviewing the predicted risk scores. Although the specific 1-to-20 value of the algorithm’s risk score likely helped workers allocate their attention more effectively, we cannot definitively rule out benefits from a general increase in the salience of family history – or other characteristics – upon seeing a listed score, regardless of its magnitude. An important topic for future research is understanding the impacts of how information is presented to users in similar settings.
@@
To what extent do human decision-makers contribute their independent expertise to improve upon algorithms? As machine learning decision supports become more widespread, so do questions about human experts’ contributions to decisions, and when to automate. In this study, outcomes improved when workers were given access to an algorithm. Workers, however, did not target investigations toward children with higher algorithm risk scores, but instead used complementary information to target more effectively on other dimensions. Given how well the algorithm risk score predicts relevant outcomes (Appendix Figure A.3), it is reasonable to ask: Could outcomes have improved even more if workers had based decisions more closely on the algorithm risk score? And would there have been implications for equity?
@@ 6.1 Counterfactual 1: Deferring Entirely to an Algorithm
The first counterfactual exercise estimates the impact of fully deferring screening decisions to the algorithm using a bounding approach. The experiment did not include an algorithm-only treatment arm due to legal constraints: the statutory requirements for a screen-in do not always correspond to high algorithm-predicted risk. By design, the algorithm provided a prediction of child risk, rather than an explicit recommendation of whether to screen in. For the counterfactual, we assume the algorithm would prioritize screen-ins for children with the highest predicted risk. The counterfactual exercise operates as follows:
@@
Results of this counterfactual exercise are plotted in Figure 5. We compare child harm under a best case for our algorithm-only scenario (dashed orange line) with outcomes observed under human-only (dashed light blue line) and human-plus-algorithm decision-making (solid blue line with confidence band). The horizontal axis varies the hypothetical effect of a screen-in for children that humans screened out but the algorithm would have screened in. Given the favorable bounding assumptions, the algorithm-only counterfactual mechanically outperforms the human-only scenario. We find that child outcomes under a human-plus-algorithm scenario are better than, though statistically indistinguishable from, this favorable algorithm-only rule under realistic screen-in treatment effects (<1<1 SD), and are also comparable under larger hypothetical treatment effects. [^19] To benchmark these magnitudes, the difference in harm between children placed in foster care and those who are not is 0.5 standard deviations of the harm index. This result highlights the limits to fully automating complex social decisions with an algorithm, and provides evidence in favor of human-algorithm complementarity in our setting.
This methodology allows the researcher to estimate bounded effects of alternative screen-in decision rules. For example, appendix Figure A.16 presents outcomes under an “oracle” screen-in decision rule where children are ranked by ex-post realized harm. According to this decision rule, the 30 percent of children with the highest ex-post realized harm index are selected to be screened in. This rule is impractical, as potential future harm is inherently uncertain at the time of decision, so the estimate represents a best-case scenario for potential gains in child wellbeing. Appendix Figure A.16 illustrates that this unrealistic “oracle” bound outperforms both humans and the human-plus-algorithm condition if one assumes large benefits from screening in (greater than a 0.5 SD reduction in harm). This finding, however, relies on the strong assumption that the two-thirds of families screened in by workers – but who would have been screened out by the oracle – are not adversely affected. Nevertheless, this oracle exercise highlights that algorithms could in principle outperform the intervention arm of our trial, perhaps by leveraging richer data sources or being better optimized for human use (Ludwig and Mullainathan, 2024).
@@ 6.2 Counterfactual 2: Marginal Increased Reliance on Algorithm
In contrast to the first counterfactual exercise, which deferred decisions entirely to the algorithm, the second counterfactual exercise examines whether social workers should have relied marginally more on the algorithm’s risk score to make screen-in decisions. Increased marginal reliance corresponds to screening in additional children that the algorithm predicted as high risk. A real-life analog of this would be a mandatory screen-in protocol for families scoring above a high-risk threshold as was implemented in Pittsburgh, PA (Lacey et al., 2024).
@@
Figure 6 shows no significant anticipated benefits from deferring marginally more to the algorithm score, relative to observed human-plus-algorithm decisions. Among children not investigated by CPS, an algorithm-only decision-maker would first prioritize screening in those with the highest risk scores, on the right side of the figure. In absence of algorithmic decision support for children with a score of 20 who were not investigated, mean harm levels were almost 0.6 SD above the mean (light blue line with circles). Social workers were missing out on a subset of high-scoring children in the control group who accrued significant harm, such that a mandatory screen-in protocol for these children might have readily improved outcomes. In contrast, when workers had access to algorithmic information (dark blue line with diamonds), screened-out children scoring a 20 only had 0.1 SD greater harm than the control group mean. With algorithm access, the reduction in harm for screened-out children on referrals scoring a 20 is statistically significant at the 90 percent confidence level ($\beta=-0.46,p=0.070$). Hence, much of the capacity for harm reduction is already realized by social workers using the algorithm freely, such that further gains from a screen-in mandate would be comparatively modest, especially relative to the costs of additional investigations that would result from such a policy.
Figure 6 confirms that although workers did not change their screen-in rates for children with the highest algorithm risk scores when using the tool (Figure A.12), workers did screen in a set of ex-post more vulnerable children. Furthermore, in situations where workers “disagreed” with the algorithm (high algorithm score, worker screen-out), harm for high-score, screened out children was comparable with the control group harm mean (horizontal red line). Therefore, ex-post partial automation of investigation decisions using algorithm predictions alone is likely to be of limited benefit in our setting. Future work with larger study samples could consider partial automation with optimal triage, as proposed by Raghu et al. (2019).
@@ 6.3 Counterfactual 3: Algorithms and Disparities without Human Oversight
Finally, we estimate with minimal assumptions how an algorithm-only decision rule would have impacted screen-in equity. This is possible through the counterfactual algorithm screen-in decision rule described in Section 6.1, which only assumes no change in the overall screen-in rate and that the algorithm would have prioritized screen-ins for children with the highest predicted risk. Figure 7 shows CPS surveillance (screen-in) disparities for relevant demographic subgroups under three decision regimes: human-only, human-plus-algorithm, and algorithm-only.
@@
Notes: This figure shows point estimates and 90 percent confidence intervals for disparities in screen-in rates by demographic category. Human-only and human-plus-algorithm disparities are directly observed in the data. Algorithm-only disparities are inferred with minimal assumptions: control group observations are sorted by algorithm risk score, and mandatory screen-ins are imposed above a score threshold that keeps overall screen-in rates constant. The vertical axis shows the disparity in screen-in rates. A zero point estimate represents zero disparity in screen-ins and is illustrated using a horizontal dashed black line. For example, in the first panel (“Black vs. Other”), screen-in rates were 20 percentage points greater for Black children compared to other children for decisions where humans did not have access to the algorithm (“Human Only”; light blue triangle on the left). Low income refers to children whose families are receiving SNAP benefits.
We find that human-plus-algorithm decisions improved equity not only compared to the status quo of human experts alone, but also compared to counterfactual screen-ins from the algorithm without human oversight. Relative to an algorithm-only decision rule, human oversight reduced disparities in CPS investigations by race, ethnicity, gender, and socioeconomic status: the dark blue squares in Figure 7 (human with algorithm) are always closer to zero than the orange diamonds (algorithm only), though point estimates are in some cases imprecise. Second, the algorithm on its own would have investigated significantly more Hispanic and low-income children relative to the status quo of human experts on their own (light blue triangles). Increased surveillance under the algorithm-only counterfactual is not associated with a greater reduction in harm for these subgroups relative to observed human-plus-algorithm decisions (Appendix Figure A.17), suggesting that the algorithm-only disparities may not be warranted. In short, the improvements in equity observed during the trial would not have been fully realized – and for some groups may have worsened – without human oversight of the algorithm. Providing algorithm support to humans leads to more equitable levels of surveillance by CPS, relative to the algorithm on its own.